OpenStudy (anonymous):

translate to a system of equations and solve susan is 10 years older than matthew. in 4 years susan will be twice as old asmatthew. how old are they now??????

OpenStudy (anonymous):

x+10=y (y+4)=2(x+4)

OpenStudy (anonymous):

ok so how old are they now

OpenStudy (anonymous):

can you solve for x and y

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

You can use one of many method to solve this 1) elimination 2) Substituion 3) Matrix

OpenStudy (anonymous):

x-y=-10

OpenStudy (anonymous):

i dont think it mattrs i can use watever works best

OpenStudy (anonymous):

Ok, In this case I think substitution is easiest x=y-10 let plug in the other equation (y+4)=2(x+4) y+4=2x+8 y-4=2x .5y-2=x ** remember x=y-10 .5y-2=y-10

OpenStudy (anonymous):

ok so far i get it but i dont think that i did it right will u postwat u got